OpenStudy (anonymous):

Given x2 + 8x - 20 = 0. Which expression below would be a correct step if solving by completing the square? A. (x + 16)2 = 36 B. (x + 4)2 = 20 C. (x + 4)2 = 36 D. (x + 16)2 = 20

jimthompson5910 (jim_thompson5910):

Work your way backwards So say we want to test if choice D is the answer (x + 16)^2 = 20 x^2 + 32x + 256 = 20 x^2 + 32x + 256 - 20 = 0 x^2 + 32x + 236 = 0 This is NOT the same as x^2 + 8x - 20 = 0, so choice D can't be the answer.

OpenStudy (anonymous):

its B right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(x + 4)^2 = 20 x^2 + 8x + 16 = 20 x^2 + 8x + 16 - 20 = 0 x^2 + 8x - 4 = 0 which is NOT the same as x^2 + 8x - 20 = 0

OpenStudy (anonymous):

ok got it..C thanks!!

jimthompson5910 (jim_thompson5910):

you got it